# From Discrete to Continuous Binary Best-Response Dynamics: Discrete Fluctuations Almost Surely Vanish with Population Size

### v2

@@ IV The link to the semicontinuous dynamics
The above dynamics are indeed the mean dynamics of the discrete population dynamics (3) [19]. The mean dynamic considers the decision-making process as a continuous process with an infinitely large population and $\dot{x}_{p}$ describes the rate of change in the population proportion of strategy-$\mathtt{A}$ players of subpopulation $p$. This equals the expected inflow, that is the rate of change to strategy $\mathtt{A}$ minus the expected outflow, that is the rate of change to strategy $\mathtt{B}$:
$\dot{x}_{p}\in F_{p,\mathtt{A}}(\bm{x})-F_{p,\mathtt{B}}(\bm{x}),$ (6)
where $F_{p,\mathtt{A}},F_{p,\mathtt{B}}:\bm{\mathcal{X}}_{ss}\rightarrow 2^{[0,\rho_{p}]}$. If the population proportion of strategy-$\mathtt{A}$ players is higher (resp. lower) than the threshold of anticoordinating (resp. coordinating) subpopulation $p$, then the inflow will be zero, and the outflow equals the current proportion of strategy-$\mathtt{A}$ players of subpopulation $p$, i.e., $\dot{x}_{p}=-x_{p}$. On the contrary, if the population proportion of strategy-$\mathtt{A}$ players is lower (resp. higher) than the threshold of anticoordinating (resp. coordinating) subpopulation $p$, then the outflow will be zero and the inflow equals the current proportion of strategy $\mathtt{B}$ players of subpopulation $p$, i.e., $\dot{x}_{p}=\rho_{p}-x_{p}$. When the population proportion of strategy-$\mathtt{A}$ players equals the threshold of (anticoordinating or coordinating) subpopulation $p$, both strategies $\mathtt{A}$ and $\mathtt{B}$ are a legitimate choice should there be no tie-breaker, and some may choose strategy $\mathtt{A}$ while others may choose strategy $\mathtt{B}$. Therefore, the rate will be bounded by the above two cases, i.e., $-x_{p}$ and $\rho_{p}-x_{p}$, resulting in the interval $[-x_{p},\rho_{p}-x_{p}]$. If there is a tie-breaker, such as the set-up in this paper, then one of the strategies is preferred, resulting in one of the above two cases, which results in a selection of the defined differential inclusion. A selection from $\bm{\mathcal{V}}(\bm{x})$ is a singleton map $\bm{\nu}(\bm{x})\in\mathbb{R}^{n}$ which satisfies $\bm{\nu}(\bm{x})\in\bm{\mathcal{V}}(\bm{x})$ for all $\bm{x}$ in $\mathbb{R}^{n}$ [41]. That is why the mean dynamics of the discrete population dynamics (3) is sometimes considered as a selection from (5) [37]. It can be easily checked that the differential inclusion (5) satisfies the conditions of good upper semicontinuity.
@@
$\dot{\bm{x}}\in\bm{\mathrm{Br}}(\bm{x})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\{}\bm{x}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\}}$
where $\bm{\mathrm{Br}}:\bm{\mathcal{X}}_{ss}\to 2^{\bm{\mathcal{X}}_{ss}}$ is the best-response correspondence where $\mathrm{Br}_{p}(\bm{x})$ returns that population proportion of strategy-$\mathtt{A}$ players of subpopulation $p$ maximizing their utility defined in the game-theoretic context of the linear threshold models (see Remark 1). In the set-up of this paper, $\mathrm{Br}_{p}(\bm{x})=\rho_{p}(2-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s^{*}}(p,\bm{x}))$.
@@
The collection of $\langle\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle_{k}\rangle_{\mathsf{N}\in\mathcal{N}}$ is a GSAP for (5).
@@
The set-up in (1), (2), and (3) is based on a tie-breaking rule [42] where the agents prefer strategy $\mathtt{A}$ when the population proportion of $\mathtt{A}$-players equals the agent’s threshold, i.e., $x^{\mathsf{N}}=\tau(i)$ for every agent $i$. However, for other tie-breaking rules, such as when strategy $\mathtt{B}$ or the current strategy of the agent is preferred, Lemma 1 remains valid. The reason is that a different tie-breaker will only change $s^{*}(\cdot,\cdot)$ and consequently the selection from (5) at $x=\tau(i)$. Therefore, investigating the asymptotic behaviour of (5) does reveal that of the discrete population dynamics (3) with such tie-breaking rules as population size approaches infinity.
V The analysis of the semicontinuous dynamics
To analyze the semicontinuous dynamics, the equilibrium points of (5) should be characterized. The following subsection provides an intuition for how analyzing the evolution of $x(t)$ in Definition 6 as the population proportion of strategy-$\mathtt{A}$ players helps that of semicontinuous population dynamics.
@@ V-A Intuition and example
Consider a mixed population of one anticoordinating and two coordinating subpopulations. The thresholds and the population proportions of the subpopulations are respectively as follows: $({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tau_{1}},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tau_{2}^{\prime}},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tau_{1}^{\prime}})=({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0.85},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0.75},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0.35})$ and $({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{1}},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\rho_{2}^{\prime}},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\rho_{1}^{\prime}})=({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0.6},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0.3},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0.1})$. The semicontinuous population dynamics are $\dot{\bm{x}}\in\bm{\mathcal{V}}(\bm{x})$, where $\bm{\mathcal{V}}=[\mathcal{V}_{1},\mathcal{V}_{2},\mathcal{V}_{3}]^{\top}$ and
@@
$\mathcal{V}_{2}(\bm{x})=\begin{cases}\{-x_{2}\}&\text{if }x<0.75,\\ [-x_{2},0.3-x_{2}]&\text{if}\quad x=0.75,\\ \{0.3-x_{2}\}&\text{if }x>0.75,\end{cases}$ (7ab)
$\mathcal{V}_{3}(\bm{x})=\begin{cases}\{-x_{3}\}&\text{if }x<0.35,\\ [-x_{3},0.1-x_{3}]&\text{if}\quad x=0.35,\\ \{0.1-x_{3}\}&\text{if }x>0.35.\end{cases}$ (7ac)
  • Motivation. To find the equilibrium points, we should find the state populations $\bm{x}^{*}$ at which $\bm{0}\in\bm{\mathcal{V}}(\bm{x}^{*})$. The combinations of all possible cases for $\mathcal{V}_{1},\mathcal{V}_{2}$, and $\mathcal{V}_{3}$ should be considered based on the value of $x$. So if we already knew the value of $x$ at equilibrium, we would only need to investigate the in-force case in $\mathcal{V}_{1}$, $\mathcal{V}_{2}$, and $\mathcal{V}_{3}$ to find $x_{1},x_{2},x_{3}$, and in turn $\bm{x}$ at equilibrium. Hence, we obtain the dynamics $x(t)$ as follows.
@@
  • Abstract dynamics’ equilibria. At equilibrium, we have $0\in\mathcal{X}(x^{*})$. So we can find the equilibrium by investigating all of the intervals in (8). The first two intervals are equilibrium-free. However, the image of the interval $(0.35,0.75)$ under $\mathcal{X}$ is $(-0.05,0.35)$, which includes zero. Therefore, the abstract dynamics admit an equilibrium point in this interval, which turns out to be $x^{*}_{1}=0.7$ as $0\in\mathcal{X}(0.7)$. Following the same procedure, we find two additional equilibrium points $x^{*}_{2}=0.75$ and $x^{*}_{3}=0.85$ for the abstract dynamics.
  • Population dynamics’ equilibria. In view of (7a), the value $x^{*}_{1}$ equals the summation of $\rho_{1}$ and $\rho^{\prime}_{1}$, i.e., the population proportions of those subpopulations whose preferred strategy at $x^{*}_{1}$ is $\mathtt{A}$. Evaluating (7a) at $x=x^{*}_{1}$, we see that $\bm{x}^{*}_{1}=(0.6,0,0.1)$ is an equilibrium point for the population dynamics.
    The point $x^{*}_{2}$ equals the threshold of the coordinating subpopulation $2$ and the population state $\bm{x}^{*}_{2}=(0.6,0.05,0.1)$ is an equilibrium point for (7a). At $\bm{x}^{*}_{2}$, except for coordinating subpopulation $2$, members of the same subpopulations adopt the same strategy, i.e., $x^{*}_{2,1}=\rho_{1}$, $x^{*}_{2,3}=\rho^{\prime}_{1}$.
    Finally, the point $x^{*}_{3}=0.85$ equals the anticoordinating subpopulation’s threshold and $\bm{x}^{*}_{3}=(0.45,0.3,0.1)$ is also an equilibrium for the population dynamics. At $\bm{x}^{*}_{3}$, except for the anticoordinating subpopulation, members of the same subpopulations adopt the same strategy, i.e., $x^{*}_{3,2}=\rho^{\prime}_{2}$, $x^{*}_{3,3}=\rho^{\prime}_{1}$.
@@
Example 3 gives us an intuition that it is useful to focus on the evolution of the population proportion of $\mathtt{A}$-players, i.e., $x(t)$, which from now on we refer to as the abstract state as it is abstracting away the heterogeneity of the population. Define ${\eta}^{\prime}_{j}$ (resp. ${\eta}_{i}$) as the cumulative population proportion of coordinators (resp. anticoordinating), that is the population proportion of those having thresholds equal to or less (resp. greater) than $\tau^{\prime}_{j}$ (resp. $\tau_{i}$), i.e.,
@@
In words, it is assumed that the summation of the cumulative population proportion of first $k$ anticoordinating subpopulations and first $l$ coordinating subpopulations does not match any of the thresholds for the $\mathsf{p}+\mathsf{p}^{\prime}$ subpopulations. We refer the reader to Remark 5 for consequences of this assumption.
@@
Consider the semicontinuous dynamics (5). Under 1, the evolution of the population proportion of $\mathtt{A}$-players $x(t)$ is governed by
@@ V-C Equilibrium points
A clean-cut population state, is a population state $\mathbf{c}^{ij},j\in[\mathsf{p}^{\prime}],i\in[\mathsf{p}],$ defined by
@@
A ruffled anticoordinator-driven population state is a population state $\mathbf{a}^{ij},j\in[\mathsf{p}^{\prime}],i\in[\mathsf{p}],$ defined by
@@
where $j$ equals $\max\{k|\tau^{\prime}_{k}<\tau_{i}\}$. The abstract state at $\mathbf{a}^{ij}$ is denoted by $a^{ij}=\tau_{i}$.
@@
A ruffled coordinator-driven population state is a population state $\mathbf{o}^{ij},j\in[\mathsf{p}^{\prime}],i\in[\mathsf{p}]$, defined by
@@
where $i=\max\{k\in[\mathsf{p}]|\tau^{\prime}_{j}<\tau_{k}\}$. The abstract state at $\mathbf{o}^{ij}$ is denoted by $o^{ij}=\tau^{\prime}_{j}$.
@@
In Example 3, $\bm{x}^{*}_{1}$ (resp. $x^{*}_{1})$ is a clean-cut equilibrium point for (7a) (resp. (8)) and equals $\mathbf{c}^{11}$ (resp. ${c}^{11}$), and $\bm{x}^{*}_{3}$ (resp. $x^{*}_{3}$) is a ruffled anticoordinator-driven equilibrium point for (7a) (resp. (8)) and equals $\mathbf{a}^{12}$ (resp. ${a}^{12}$). Finally, $\bm{x}^{*}_{2}$ (resp. $x^{*}_{2}$) is a ruffled coordinator-driven equilibrium point for (7a) (resp. (8)) and equals $\mathbf{o}^{12}$ (resp. ${o}^{12}$). From now on, for the sake of readability, we drop the word “ruffled” and simply say “(anti)coordinator-driven.” The following lemma proves that the equilibria of the semicontinuous population and abstract dynamics are one of these three population states.
@@
According to Lemma 2, for every equilibrium of the semicontinuous population dynamics, e.g., $\mathbf{c}^{ij}$, there is exactly one associated equilibrium in the abstract dynamics, i.e., $c^{ij}$.

### v3

@@ IV The link to the continuous-time dynamics
The above dynamics are indeed the mean dynamics of the discrete population dynamics (3) [20]. The mean dynamic considers the decision-making process as a continuous process with an infinitely large population, and $\dot{x}_{p}$ describes the rate of change in the population proportion of strategy-$\mathtt{A}$ players in subpopulation $p$. This equals the expected inflow,$F_{p,\mathtt{A}}$, that is, the rate of change to strategy $\mathtt{A}$ minus the expected outflow, $F_{p,\mathtt{B}}$, that is, the rate of change to strategy $\mathtt{B}$:
$\dot{x}_{p}\in{\{}F_{p,\mathtt{A}}(\bm{x})-F_{p,\mathtt{B}}(\bm{x}){\}},$ (8)
where $F_{p,\mathtt{A}},F_{p,\mathtt{B}}:\bm{\mathcal{X}}_{ss}\rightarrow{[0,\rho_{p}]}$. If the population proportion of strategy-$\mathtt{A}$ players is greater (resp. less) than the threshold of anticoordinating (resp. coordinating) subpopulation $p$, then the inflow will be zero, and the outflow equals the current proportion of strategy-$\mathtt{A}$ players of subpopulation $p$, i.e., $\dot{x}_{p}=-x_{p}$. On the contrary, if the population proportion of strategy-$\mathtt{A}$ players is less (resp. greater) than the threshold of anticoordinating (resp. coordinating) subpopulation $p$, then the outflow will be zero and the inflow equals the current proportion of strategy $\mathtt{B}$ players of subpopulation $p$, i.e., $\dot{x}_{p}=\rho_{p}-x_{p}$. When the population proportion of strategy-$\mathtt{A}$ players equals the threshold of (anticoordinating or coordinating) subpopulation $p$, both strategies $\mathtt{A}$ and $\mathtt{B}$ are a legitimate choice should there be no tie-breaker, and some may choose strategy $\mathtt{A}$ while others may choose strategy $\mathtt{B}$. Therefore, the rate will be bounded by the above two cases, i.e., $-x_{p}$ and $\rho_{p}-x_{p}$, resulting in the interval $[-x_{p},\rho_{p}-x_{p}]$. If there is a tie-breaker, such as the set-up in this paper, then one of the strategies is preferred, resulting in one of the above two cases, which results in a selection from the defined differential inclusion. A selection from $\bm{\mathcal{V}}(\bm{x})$ is a singleton map $\bm{\nu}(\bm{x})\in\mathbb{R}^{n}$ which satisfies $\bm{\nu}(\bm{x})\in\bm{\mathcal{V}}(\bm{x})$ for all $\bm{x}$ in $\mathbb{R}^{n}$ [42]. That is why the mean dynamics of the discrete population dynamics (3) is sometimes considered as a selection from (7) [38]. It can be easily checked that the differential inclusion (7) satisfies the conditions of good upper semicontinuity.
@@
$\dot{\bm{x}}\in\bm{\mathrm{Br}}(\bm{x})-\bm{x}$
where $\bm{\mathrm{Br}}:\bm{\mathcal{X}}_{ss}{\rightrightarrows}{\bm{\mathcal{X}}_{ss}}$ is the best-response correspondence where $\mathrm{Br}_{p}(\bm{x})$ returns that population proportion of strategy-$\mathtt{A}$ players of subpopulation $p$ maximizing their utility defined in the game-theoretic context of the linear threshold models (see Remark 1).
@@
The collection of $\langle\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle{{{}_{k=0}^{\infty}}}\rangle_{\mathsf{N}\in\mathcal{N}}$ is a GSAP for (7).
@@
The set-up in (1), (2), and (3) is based on a tie-breaking rule [43] where the agents prefer strategy $\mathtt{A}$ when the population proportion of $\mathtt{A}$-players, excluding themselves, equals the agent’s threshold, i.e., $x^{\mathsf{N}}_{-i}=\tau(i)$ for every agent $i$. However, for other tie-breaking rules, such as when strategy $\mathtt{B}$ is preferred, or the agents uniformly randomly choose either of the strategies, Lemma 1 remains valid–A1. The reason is that a different tie-breaker will only change $s^{*}(\cdot,\cdot,\cdot)$ in Definition 1 and consequently the selection from (7) when $x=\tau_{p}$ where $p$ is the type of agent $i$. Accordingly, the dynamical behavior of the resulting selection is already captured by (7). Therefore, investigating the asymptotic behavior of (7) does reveal that of the discrete population dynamics (3) with such tie-breaking rules as population size approaches infinity.
V The analysis of the continuous-time dynamics
To analyze the continuous-time dynamics, the equilibrium points of (7) should be characterized. The following subsection provides an intuition for how analyzing the evolution of $x(t)$ in Definition 6, i.e., the population proportion of strategy-$\mathtt{A}$ players, helps that of continuous-time population dynamics.
@@ V-A Intuition and example
Consider a mixed population of one anticoordinating and two coordinating subpopulations. The thresholds and the population proportions of the subpopulations are respectively as follows: $({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tau_{1}},{\tau_{2}^{\prime}},{\tau_{1}^{\prime}})=({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0.85},{0.75},{0.35})$ and $({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{1}},{\rho_{2}^{\prime}},{\rho_{1}^{\prime}})=({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0.6},{0.3},{0.1})$. The continuous-time population dynamics are $\dot{\bm{x}}\in\bm{\mathcal{V}}(\bm{x})$, where $\bm{\mathcal{V}}=[\mathcal{V}_{1},\mathcal{V}_{2},\mathcal{V}_{3}]^{\top}$ and
@@
$\mathcal{V}_{2}(\bm{x})=\begin{cases}\{-x_{2}\}&\text{if }x<0.75,\\ [-x_{2},0.3-x_{2}]&\text{if}\quad x=0.75,\\ \{0.3-x_{2}\}&\text{if }x>0.75,\end{cases}$ (9ab)
$\mathcal{V}_{3}(\bm{x})=\begin{cases}\{-x_{3}\}&\text{if }x<0.35,\\ [-x_{3},0.1-x_{3}]&\text{if}\quad x=0.35,\\ \{0.1-x_{3}\}&\text{if }x>0.35.\end{cases}$ (9ac)
  • Motivation. To find the equilibrium points, we should find the state populations $\bm{x}^{*}$ at which $\bm{0}\in\bm{\mathcal{V}}(\bm{x}^{*})$. The combinations of all possible cases for $\mathcal{V}_{1},\mathcal{V}_{2}$, and $\mathcal{V}_{3}$ should be considered based on the value of $x$. So if we already knew the value of $x$ at equilibrium, we would only need to investigate the active case in $\mathcal{V}_{1}$, $\mathcal{V}_{2}$, and $\mathcal{V}_{3}$ to find $x_{1},x_{2},x_{3}$, and in turn $\bm{x}$ at equilibrium. Hence, we obtain the dynamics $x(t)$ as follows.
@@
  • Abstract dynamics’ equilibria. At equilibrium, we have $0\in\mathcal{X}(x^{*})$. So we can find the equilibria by investigating all intervals in (10). The first two intervals are equilibrium-free. However, the image of the interval $(0.35,0.75)$ under $\mathcal{X}$ is $(-0.05,0.35)$, which includes zero. Therefore, the abstract dynamics admit an equilibrium point in this interval, which turns out to be $x^{*}_{1}=0.7$ as $0\in\mathcal{X}(0.7)$. Following the same procedure, we find two additional equilibrium points $x^{*}_{2}=0.75$ and $x^{*}_{3}=0.85$ for the abstract dynamics.
  • Population dynamics’ equilibria. In view of (9a), the value $x^{*}_{1}$ equals the sum of $\rho_{1}$ and $\rho^{\prime}_{1}$, i.e., the population proportions of those subpopulations whose preferred strategy at $x^{*}_{1}$ is $\mathtt{A}$. Evaluating (9a) at $x=x^{*}_{1}$, we see that $\bm{x}^{*}_{1}=(0.6,0,0.1)$ is an equilibrium point for the population dynamics. The point $x^{*}_{2}$ equals the threshold of the second coordinating subpopulation, and the population state $\bm{x}^{*}_{2}=(0.6,0.05,0.1)$ is an equilibrium point for (9a). At $\bm{x}^{*}_{2}$, except for the second coordinating subpopulation, members of the same subpopulations adopt the same strategies, i.e., $x^{*}_{2,1}=\rho_{1}$, $x^{*}_{2,3}=\rho^{\prime}_{1}$. Finally, the point $x^{*}_{3}=0.85$ equals the anticoordinating subpopulation’s threshold, and $\bm{x}^{*}_{3}=(0.45,0.3,0.1)$ is an equilibrium for the population dynamics. At $\bm{x}^{*}_{3}$, except for the anticoordinators, members of the same subpopulations adopt the same strategies, i.e., $x^{*}_{3,2}=\rho^{\prime}_{2}$, $x^{*}_{3,3}=\rho^{\prime}_{1}$.
@@
Example 3 gives us an intuition that it is useful to focus on the evolution of the population proportion of $\mathtt{A}$-players, i.e., $x(t)$, which from now on we refer to as the abstract state as it is abstracting away the heterogeneity of the population. Define ${\eta}^{\prime}_{j}$ (resp. ${\eta}_{i}$) as the cumulative population proportion of coordinating (resp. anticoordinating) subpopulations, that is the population proportion of those having thresholds equal to or less (resp. greater) than $\tau^{\prime}_{j}$ (resp. $\tau_{i}$), i.e.,
@@
In words, it is assumed that the sum of the cumulative population proportion of first $k$ anticoordinating subpopulations and first $l$ coordinating subpopulations does not match any of the thresholds for the $\mathsf{p}+\mathsf{p}^{\prime}$ subpopulations. We refer the reader to Remark 4 for consequences of this assumption.
@@
Consider the continuous-time dynamics (7). Under 1, the evolution of the population proportion of $\mathtt{A}$-players $x(t)$ is governed by
@@ V-C Equilibrium points
A clean-cut population state, is a population state $\mathbf{c}^{ij},i\in[\mathsf{p}],j\in[\mathsf{p}^{\prime}],$ defined by
@@
A ruffled anticoordinator-driven population state is a population state $\mathbf{a}^{ij},i\in[\mathsf{p}],j\in[\mathsf{p}^{\prime}],$ defined by
@@
where $j$ equals $\max\{k\in[\mathsf{p}^{\prime}]\cup\{0\}|\tau^{\prime}_{k}<\tau_{i}\}$. The abstract state at $\mathbf{a}^{ij}$ is denoted by $a^{ij}=\tau_{i}$. For $j=0$, no coordinators play $\mathtt{A}$.
@@
A ruffled coordinator-driven population state is a population state $\mathbf{o}^{ij},i\in[\mathsf{p}],j\in[\mathsf{p}^{\prime}]$, defined by
@@
where $i=\max\{k\in[\mathsf{p}]\cup\{0\}|\tau^{\prime}_{j}<\tau_{k}\}$. The abstract state at $\mathbf{o}^{ij}$ is denoted by $o^{ij}=\tau^{\prime}_{j}$. For $i=0$, no anticoordinators play $\mathtt{A}$.
@@
In Example 3, $\bm{x}^{*}_{1}$ (resp. $x^{*}_{1})$ is a clean-cut equilibrium point for (9a) (resp. (10)) and equals $\mathbf{c}^{11}$ (resp. ${c}^{11}$), and $\bm{x}^{*}_{3}$ (resp. $x^{*}_{3}$) is a ruffled anticoordinator-driven equilibrium point for (9a) (resp. (10)) and equals $\mathbf{a}^{12}$ (resp. ${a}^{12}$). Finally, $\bm{x}^{*}_{2}$ (resp. $x^{*}_{2}$) is a ruffled coordinator-driven equilibrium point for (9a) (resp. (10)) and equals $\mathbf{o}^{12}$ (resp. ${o}^{12}$). From now on, for the sake of readability, we drop the word “ruffled” and simply say “(anti)coordinator-driven.” The following lemma proves that the equilibria of the continuous-time population and abstract dynamics are one of these three population states.
@@
According to Lemma 2, for every equilibrium of the continuous-time population dynamics, e.g., $\mathbf{c}^{ij}$, there is exactly one associated abstract equilibrium point, i.e., $c^{ij}$.